The operation manager at a tire manufacturing company believes that the mean mileage of a tire is 49,36049,360 miles, with a variance of 6,568,9696,568,969.

What is the probability that the sample mean would differ from the population mean by less than 231231 miles in a sample of 247247 tires if the manager is correct? Round your answer to four decimal places.

CAN YOU PLEASE SHOW ME HOW TO DO THIS IN CALCULATOR TI-83 PLUS SPECFICALLY
Answer #1

SD = sqrt(6568969/247) = 163.0798

Lower = 49360-231 = 49129

Upper = 49360+231 = 49591

2ND+VARS --> normalcdf

normalcdi lower: 49129 upper: 49591 .: 49360 o: 163.0798 TAT PLOT TBLSET 32 FORMAT CALC WINDOW TABLE GRAPH ZOOM TRACE

TEXAS INSTRUMENTS normalcdf (49129 .8433661877

Ans : 0.8434